OpenStudy (anonymous):

Help!! @kelliegirl33 @UsArmy3947 Pre-Algebra

OpenStudy (anonymous):

8n + 2 = 3n .....please rewrite this problem

OpenStudy (anonymous):

is that supposed to be 3n + 13 ??

OpenStudy (anonymous):

oh sorry i was ordering a pizza... The equation is supposed to say 8n+2=3n-13

OpenStudy (anonymous):

8n + 2 = 3n - 13 (subtract 3n from both sides) 8n - 3n + 2 = -13 (subtract 2 from both sides) 8n - 3n = -13 - 2 (combine like terms) 5n = -15 (divide by 5) n = -3 check... 8(-3) + 2 = 3(-3) - 13 -24 + 2 = -9 - 13 -22 = - 22 (correct) n = -3 pizza sound good right now...lol

OpenStudy (anonymous):

here is a little tip....whenever you move something across the equal sign, just change the sign. 2x + 4 = 10 -- you want to move the 4 across the equal sign...just move it and change the sign. 2x = 10 - 4 oh..another problem....let me check it out...oh..it disappeared

OpenStudy (anonymous):

its back....hold on...

OpenStudy (anonymous):

5(x + 3) = ......where did it go....I started to write it down

OpenStudy (anonymous):

Which of the following is a simplified form of the expression five (x plus three) minus two (x minus y)? seven x minus two y plus fifteen three x plus two y plus fifteen seven x plus two y plus fifteen three x plus thirteen y

OpenStudy (anonymous):

5(x + 3) - 2(x - y) -- distribute through the parenthesis 5x + 15 - 2x + 2y -- combine like terms 3x + 2y + 15

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

wow!! Thanks to you helping me I got a 90.62% on my homework!! :) Thanks so Much!!

OpenStudy (anonymous):

cool....good grade. Glad I could help :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

:)